# Token Prepending: A Training-Free Approach for Eliciting Better Sentence Embeddings from LLMs

### v1

@@ Abstract
Extracting sentence embeddings from large language models (LLMs) is a promising direction, as LLMs have demonstrated stronger semantic understanding capabilities. Previous studies typically focus on prompt engineering to elicit sentence embeddings from LLMs by prompting the model to encode sentence information into the embedding of the last token. However, LLMs are mostly decoder-only models with causal attention and the earlier tokens in the sentence cannot attend to the latter tokens, resulting in biased encoding of sentence information and cascading effects on the final decoded token. To this end, we propose a novel Token Prepending (TP) technique that prepends each layer’s decoded sentence embedding to the beginning of the sentence in the next layer’s input, allowing earlier tokens to attend to the complete sentence information under the causal attention mechanism. The proposed TP technique is a plug-and-play and training-free technique, which means it can be seamlessly integrated with various prompt-based sentence embedding methods and autoregressive LLMs. Extensive experiments on various Semantic Textual Similarity (STS) tasks and downstream classification tasks demonstrate that our proposed TP technique can significantly improve the performance of existing prompt-based sentence embedding methods across different LLMs, while incurring negligible additional inference cost.
@@
Yuchen Fu [^1] Equal Contribution.  Zifeng Cheng [^1] Equal Contribution.  Zhiwei Jiang [^2] Corresponding Author.  Zhonghui Wang Yafeng YinZhengliang LiQing Gu State Key Laboratory for Novel Software Technology, Nanjing University, China {yuchenfu,chengzf}@smail.nju.edu.cn, jzw@nju.edu.cn zhonghuiwang@smail.nju.edu.cn, yafeng@nju.edu.cn lzl@smail.nju.edu.cn, guq@nju.edu.cn
@@ 1 Introduction
In this paper, we propose a simple yet effective technique called Token Prepending (TP). As shown in Figure 1(b), our core idea is to prepends each layer’s decoded sentence embedding to the beginning of the sentence in the next layer’s input, allowing earlier tokens to attend to the complete sentence information under the causal attention mechanism. Notably, the TP technique is entirely training-free, as it introduces no additional learnable parameters. Specifically, although TP is applicable to all layers, we find that it is not necessary to perform the TP operation across all layers. Instead, performing this operation only in the early layers of the model yields better performance. Therefore, we discontinue the TP operation after several early layers and revert to standard forward propagation. Additionally, considering that the final layer of LLMs is primarily used for token generation and contains less semantic information, we propose an early-exit strategy that outputs embeddings from intermediate layers, rather than the final layer, to serve as sentence embeddings.
@@ 2 Related Work
Extracting Sentence Embeddings from LLMsExisting methods on extracting sentence embeddings from LLMs mainly focus on designing prompts to improve sentence embeddings. PromptEOL Jiang et al. (2023) demonstrates the potential of LLMs in generating sentence embeddings by leveraging prompt engineering. Echo embeddings Springer et al. (2024) repeats the input twice within the context and extracts embeddings from the second occurrence, allowing early token embeddings to encode information about subsequent tokens. MetaEOL Lei et al. (2024) designs meta-task prompts via ChatGPT-4 to guide LLMs to consider sentence representations from multiple perspectives. Pretended CoT Zhang et al. (2024) uses CoT to inspire the model to output better embeddings. Knowledge Enhancement Zhang et al. (2024) provides explicit guidance to the model by conveying human experience in text summarization through prompts. In this paper, we propose a plug-and-play technique TP to improve the various prompt-based methods with negligible additional inference cost.
@@ 3 Preliminary
This sentence: “[Text]" means in one word: “
@@
Formally, given the input $T=[t_{1},...,t_{n}]$ wrapped in a template, we first obtain the embeddings $\textbf{h}^{0}=[h^{0}_{1},\cdots,h^{0}_{n}]$ through the embedding layer, and then pass them into the $L$ Transformer layers of LLMs. Finally, we use the last layer’s hidden state for the sentence embedding token $\textbf{h}^{L}_{n}$ as the output sentence embedding. Specifically,
@@ 4.2.2 Intermediate Token Prepending
where $k\in[2,L]$ denotes the ending layer for the intermediate token prepending and $i^{*}$ is the position index of the <PST> token.
@@
Table 1: Results on STS tasks (Spearman correlation scaled by 100x) using LLaMA2-7B as backbone. The Time column refers to the ratio of inference time for various prompt methods relative to PromptEOL on the STS-B test dataset, ensuring the same output layer and a batch size of 16.
@@ 4.3 Early-Exit from Intermediate Layers
Recent studies Liu et al. (2024b); Jin et al. (2024) demonstrate that each layer of LLMs plays a different role, and the embeddings from the last layer are primarily used for prediction and contain weaker semantic information. Thus, we propose the early-exit strategy, which uses embeddings from intermediate layers instead of the last layer to serve as sentence embeddings. We use the validation set to determine which layer of embedding to use, and the overhead of this process is light. Another advantage of the early-exit strategy is that we can obtain sentence embeddings more quickly during the testing phase.
@@ 5.2 Baselines
We combine our method with some baselines to demonstrate effectiveness. BERT avg Devlin et al. (2019), ST5-End avg Ni et al. (2022a), and LLaMA2 avg Touvron et al. (2023) average token embeddings to obtain sentence embeddings using different backbones. LLaMA2 echo Springer et al. (2024) utlizes the strategy of repetition to obtain sentence embeddings. BERT prompt Jiang et al. (2023) proposes a simple and effective prompt to extract sentence embeddings from BERT. PromptEOL Jiang et al. (2023) first proposes a simple and effective prompt to extract sentence embeddings from LLMs. MetaEOL Lei et al. (2024) leverages a diverse set of meta-task prompts to capture multiple representations of sentences from distinct perspectives. Pretended CoT Zhang et al. (2024) uses CoT to inspire the model to extract sentence embeddings. Knowledge Zhang et al. (2024) explicitly infuses the model with human insights into text summarization.
@@
        This sentence : <PST> “[Text]" means in one word:¨         77.19
        <PST> This sentence : “[Text]" means in one word:¨         76.35
        This sentence : “<PST> [Text]" means in one word:¨         76.71
        After thinking step by step , this sentence : <PST> “[Text]" means in one word:¨         77.54
        <PST> After thinking step by step , this sentence : “[Text]" means in one word:¨         76.53
        After thinking step by step , <PST> this sentence : “[Text]" means in one word:¨         77.81
        After thinking step by step , this sentence : “<PST> [Text]" means in one word:¨         77.51
@@ 5.5 Analysis of <PST> Token
Effects of the position of <PST>We employ PromptEOL and Pretended CoT as the prompt template to examine how the placement of the <PST> token at different positions in the sentence affects performance on STS tasks, as shown in Table 3. The performance is worse when <PST> token is positioned at the beginning of the prompt. When <PST> token is placed near the text, performance fluctuation is smaller. The optimal position of <PST> token varies depending on the prompt, typically positioning it close to the text. To avoid the additional overhead of searching the position, we simply place <PST> token after the colon for all prompts.
@@
Effectiveness of retaining the <PST> token before and after intermediate TP We ablate the <PST> token before and after intermediate token prepending to show the effectiveness of retaining the <PST> token. Ablating the <PST> token before the intermediate token prepending is equivalent to ablating the initial token prepending and directly performing intermediate token prepending.
@@ 5.6 Analysis of Layer Scope for TP
As illustrated in Figure 4(a), performance declines if intermediate token prepending does not begin at the second layer. This is because the <PST> token is randomly initialized and lacks semantic information. Thus, we need to replace it with semantically meaningful tokens at the early layers of the LLMs to mitigate this issue.
@@ 5.8 Transfer Learning Tasks
The results results for the transfer tasks, shown in Table 5, demonstrate that our method consistently outperforms all baselines, with improvements in 20 out of 21 cases across all datasets. This indicates that TP cultivates generalized sentence embeddings that perform outstandingly across various tasks. Pretended CoT and Knowledge Enhancement do not surpass the performance of PromptEOL, indicating that they are not consistently effective in enhancing performance across tasks.
@@ 6 Conclusion
In this paper, we introduce Token Prepending technique, a plug-and-play approach for deriving high-quality sentence embeddings from autogressive LLMs without requiring any training and data. By intervening in the inputs to Transformer layers, TP enhances the ability of autoregressive LLMs to capture backward dependencies. Moreover, TP involves simply prepending a single token to the sentence, which adds negligible inference cost and can seamlessly integrates with prompt-based methods. Our extensive experiments demonstrate that TP technique can effectively and generally elicit sentence embeddings across a range of LLMs with varying architectures and parameter sizes, achieving outstanding performance on both STS datasets and transfer learning tasks. We find that starting TP from the first layer yields optimal results, and the best stopping point is typically around the 7th or 8th layer for LLMs with about 7B parameters.

### v2

@@ Abstract
Extracting sentence embeddings from large language models (LLMs) is a promising direction, as LLMs have demonstrated stronger semantic understanding capabilities. Previous studies typically focus on prompt engineering to elicit sentence embeddings from LLMs by prompting the model to encode sentence information into the embedding of the last token. However, LLMs are mostly decoder-only models with causal attention and the earlier tokens in the sentence cannot attend to the latter tokens, resulting in biased encoding of sentence information and cascading effects on the final decoded token. To this end, we propose a novel Token Prepending (TP) technique that prepends each layer’s decoded sentence embedding to the beginning of the sentence in the next layer’s input, allowing earlier tokens to attend to the complete sentence information under the causal attention mechanism. The proposed TP technique is a plug-and-play and training-free technique, which means it can be seamlessly integrated with various prompt-based sentence embedding methods and autoregressive LLMs. Extensive experiments on various Semantic Textual Similarity (STS) tasks and downstream classification tasks demonstrate that our proposed TP technique can significantly improve the performance of existing prompt-based sentence embedding methods across different LLMs, while incurring negligible additional inference cost. The code are available on https://github.com/fuyuchenIfyw/token_prepending.git.
@@
Yuchen Fu11footnotemark: 1  Zifeng Cheng11footnotemark: 1  Zhiwei Jiang22footnotemark: 2  Zhonghui Wang Yafeng YinZhengliang LiQing Gu State Key Laboratory for Novel Software Technology, Nanjing University, China {yuchenfu,chengzf}@smail.nju.edu.cn, jzw@nju.edu.cn zhonghuiwang@smail.nju.edu.cn, yafeng@nju.edu.cn lzl@smail.nju.edu.cn, guq@nju.edu.cn
@@ 1 Introduction
In this paper, we propose a simple yet effective technique called Token Prepending (TP). As shown in Figure 1(b), our core idea is to prepends each layer’s decoded sentence embedding to the beginning of the sentence in the next layer’s input, allowing earlier tokens to attend to the complete sentence information under the causal attention mechanism. Notably, the TP technique is entirely training-free, as it introduces no additional learnable parameters. Specifically, although TP is applicable to all layers, we find that it is not necessary to perform the TP operation across all layers. Instead, performing this operation only in the early layers of the model yields better performance. Therefore, we discontinue the TP operation after several early layers and revert to standard forward propagation. Additionally, considering that the final layer of LLMs is primarily used for token generation and contains less semantic information Liu et al. (2024b); Jin et al. (2024b), we propose an early-exit strategy that outputs embeddings from intermediate layers, rather than the final layer, to serve as sentence embeddings.
@@ 2 Related Work
Extracting Sentence Embeddings from LLMsExisting methods on extracting sentence embeddings from LLMs mainly focus on designing prompts to improve sentence embeddings. PromptEOL Jiang et al. (2023) demonstrates the potential of LLMs in generating sentence embeddings by leveraging prompt engineering. Echo embeddings Springer et al. (2024) repeats the input twice within the context and extracts embeddings from the second occurrence, allowing early token embeddings to encode information about subsequent tokens. MetaEOL Lei et al. (2024) designs meta-task prompts via ChatGPT-4 to guide LLMs to consider sentence representations from multiple perspectives. Pretended CoT Zhang et al. (2024) uses CoT to inspire the model to output better embeddings. Knowledge Enhancement Zhang et al. (2024) provides explicit guidance to the model by conveying human experience in text summarization through prompts. CP Cheng et al. (2025) introduces an extra auxiliary prompt to elicit better sentence embedding. In this paper, we propose a plug-and-play technique TP to improve the various prompt-based methods with negligible additional inference cost.
@@ 3 Preliminary
This sentence: “[Text]” means in one word: “
@@
Formally, given the input $T=[t_{1},...,t_{n}]$ wrapped in a template, we first obtain the embeddings $\textbf{h}^{0}=[h^{0}_{1},\cdots,h^{0}_{n}]$ through the embedding layer, and then pass them into the $L$ Transformer layers of LLMs. The previous work (e.g., PromptEOL) use the last layer’s hidden state for the sentence embedding token $\textbf{h}^{L}_{n}$ as the output sentence embedding. Specifically,
@@ 4.2.2 Intermediate Token Prepending
where $f(\textbf{h})$ represents the function that operates on h. $k\in[2,L]$ denotes the ending layer for the intermediate token prepending and $i^{*}$ is the position index of the <PST> token.
@@
Table 1: Results on STS tasks (Spearman correlation scaled by 100x) using LLaMA2-7B as backbone. The Time column refers to the ratio of inference time for various prompt methods relative to PromptEOL on the STS-B test dataset, ensuring the same output layer.
@@ 4.3 Early-Exit from Intermediate Layers
Recent studies Liu et al. (2024c); Jin et al. (2024a) demonstrate that each layer of LLMs plays a different role, and the embeddings from the last layer are primarily used for prediction and contain weaker semantic information. Thus, we propose the early-exit strategy, which uses embeddings from intermediate layers instead of the last layer to serve as sentence embeddings. We use the validation set to determine which layer of embedding to use, and the overhead of this process is light. Another advantage of the early-exit strategy is that we can obtain sentence embeddings more quickly during the testing phase.
@@ 5.2 Baselines
We combine our method with some baselines to demonstrate effectiveness. BERT avg Devlin et al. (2019), ST5-Enc avg Ni et al. (2022a), and LLaMA2 avg Touvron et al. (2023) average token embeddings to obtain sentence embeddings using different backbones. LLaMA2 echo Springer et al. (2024) utlizes the strategy of repetition to obtain sentence embeddings. BERT prompt Jiang et al. (2022) proposes a simple and effective prompt to extract sentence embeddings from BERT. PromptEOL Jiang et al. (2023) first proposes a simple and effective prompt to extract sentence embeddings from LLMs. MetaEOL Lei et al. (2024) leverages a diverse set of meta-task prompts to capture multiple representations of sentences from distinct perspectives. Pretended CoT Zhang et al. (2024) uses CoT to inspire the model to extract sentence embeddings. Knowledge Zhang et al. (2024) explicitly infuses the model with human insights into text summarization.
@@
        This sentence : <PST> “[Text]” means in one word: “         77.19
        <PST> This sentence : “[Text]” means in one word: “         76.35
        This sentence : “<PST> [Text]” means in one word: “         76.71
        This sentence : “ [Text]” <PST> means in one word: “         75.54
        After thinking step by step , this sentence : <PST> “[Text]” means in one word: “         77.54
        After thinking step by step , <PST> this sentence : “[Text]” means in one word: “         77.81
        After thinking step by step , this sentence : “<PST> [Text]” means in one word: “         77.51
        After thinking step by step , this sentence : “ [Text]” <PST> means in one word: “         77.44
@@ 5.5 Analysis of <PST> Token
Effects of the position of <PST>We employ PromptEOL and Pretended CoT as the prompt template to examine how the placement of the <PST> token at different positions in the sentence affects performance on STS tasks, as shown in Table 3. The performance is worst when <PST> token is inserted right after the input text. When <PST> token is placed before the text, performance fluctuation is small. The optimal position of <PST> token varies depending on the prompt, typically positioning it close to the text. To avoid the additional overhead of searching the position, we simply place <PST> token after the colon for all prompts.
@@
Effectiveness of retaining the <PST> token before and after intermediate TP We ablate the <PST> token before and after intermediate token prepending to show the effectiveness of retaining the <PST> token. Ablating the <PST> token before the intermediate token prepending is equivalent to removing the initial token prepending and directly performing intermediate token prepending.
@@ 5.6 Analysis of Layer Scope for TP
As illustrated in Figure 4(a), performance is suboptimal if intermediate token prepending does not begin at the second layer. This is because the <PST> token is randomly initialized and lacks semantic information. Thus, we need to replace it with semantically meaningful tokens at the early layers of the LLMs to mitigate this issue.
@@ 5.8 Transfer Learning Tasks
The results for the transfer tasks, shown in Table 5, demonstrate that our method consistently outperforms all baselines, with improvements in 20 out of 21 cases across all datasets. This indicates that TP cultivates generalized sentence embeddings that perform outstandingly across various tasks. Pretended CoT and Knowledge Enhancement do not surpass the performance of PromptEOL, indicating that they are not consistently effective in enhancing performance across tasks.
@@ 6 Conclusion
In this paper, we introduce Token Prepending technique, a plug-and-play approach for deriving high-quality sentence embeddings from autogressive LLMs without requiring any training and data. By intervening in the inputs to Transformer layers, TP enhances the ability of autoregressive LLMs to capture backward dependencies. Moreover, TP involves simply prepending a single token to the sentence, which adds negligible inference cost and can seamlessly integrate with prompt-based methods. Our extensive experiments demonstrate that TP technique can effectively and generally elicit sentence embeddings across a range of LLMs with varying architectures and parameter sizes, achieving outstanding performance on both STS datasets and transfer learning tasks. We find that starting TP from the first layer yields optimal results, and the best stopping point is typically around the 7th or 8th layer for LLMs with about 7B parameters.
@@
Acknowledgments
We would like to thank the anonymous reviewers for their insightful comments. This work is supported by the National Natural Science Foundation of China under Grants Nos. 62441225, 61972192, 62172208, 61906085. This work is partially supported by Collaborative Innovation Center of Novel Software Technology and Industrialization. This work is supported by the Fundamental Research Funds for the Central Universities under Grant No.14380001.
@@ A.3 More Prompt Baseline Evaluation
Method Backbone STS12 STS13 STS14 STS15 STS16 STS-B SICK-R Avg.
Resuming at Layer 9 LLaMA2-7B 66.97 82.75 73.24 78.75 79.36 80.40 75.33 76.69
Resuming at Layer 10 LLaMA2-7B 66.76 82.91 73.26 79.11 79.40 80.30 75.58 76.76
Resuming at Layer 11 LLaMA2-7B 66.68 83.01 73.34 79.13 79.46 80.17 75.60 76.77
Resuming at Layer 16 LLaMA2-7B 66.84 83.07 74.22 79.83 79.96 80.56 75.36 77.12
Resuming at Layer 26 LLaMA2-7B 66.95 83.18 74.33 79.89 80.04 80.66 75.40 77.21
Table 16: Results on STS tasks (Spearman correlation scaled by 100x) based on PromptEOL, resuming TP at different layer.
Method Backbone STS12 STS13 STS14 STS15 STS16 STS-B SICK-R Avg.
w/o prompt LLaMA2-7B 9.13 22.25 11.04 33.09 34.92 15.94 33.73 22.87
w/o prompt + TP LLaMA2-7B 11.68 25.85 13.08 33.70 46.58 22.05 42.61 27.94
Table 17: Results on STS tasks (Spearman correlation scaled by 100x) based on PromptEOL without any prompt.
@@ A.5 Masking <PST> Token in the First Layer
A.6 Resuming TP at Different Layers
We explore the practice of using TP for a few layers, and then pause for a few layers and then resume again. We conduct experiments based on the best practices of PromptEOL and TP, applying TP at the 1st layer, stopping TP at the 8th layer, and early exiting at the 27th layer. The results are shown in the Table 16. Once resumed, TP remains active until output. We find that resuming TP at deeper layers (e.g., layer 21) yields a slight performance improvement.
A.7 TP without Prompt
We test performance of TP without any prompt. The results are shown in the Table17. Although the absence of a prompt significantly degrades performance, TP still manages to provide improvements.